# Sequence Generation Modeling for Continuous Value Prediction

### v2

@@ 4.4. Generative Architecture Analysis (RQ3)
4.5. Effect of Curriculum Learning with Embedding Mixup (RQ4)
Here, we investigate the impact of the proposed curriculum learning (CL) with embedding mixup (EM) strategy by designing several variants to perform comparisons from three perspectives.
  • To study the contribution of CLEM, we design three ablation variants. w/o CLEM uses free running mode to directly feed $\bm{E}[\hat{s}_{i}^{t-1},:]$ to generate $\hat{s}_{i}^{t-1}$. EM with TF indicates embedding mixup strategy with full teacher forcing, i.e., with fixed sampling rate $p$ = 1. CL w/o EM removes embedding mixup strategy.
  • To explore the effect of different dynamic decay strategy, we design linear decay and exponential decay instead of Eq. (7), thereby constructing two variants linear and exponential.
  • To examine the effect of the curriculum strategy and free running, we construct two model variants with fixed sampling rates, where $p$ is set to 0.5 and 0, respectively.
Table 5. Ablation study on the strategy of curriculum learning with embedding mixup.
@@ 4.5. Effect of Curriculum Learning with Embedding Mixup (RQ4)
The results in Tab. 5 indicate that both strategies in CLEM enhance performance. Specifically, comparing row (a) with row (c), we can see that CL increases XAUC by 1.656% and reduces MAE by 1.38% on the KuaiRec dataset, while comparing row (a) with row (d) shows EM achieves even greater improvements, with 4.235% increase in XAUC and 4.853% reduction in MAE. Removing all the strategies results in substantial performance degradation for the GR model, with a loss of 5.137% in XAUC and 6.884% in MAE.
The sampling rate decay coefficients significantly impact both metrics. The proposed curriculum strategy achieves a gain of 2.65% in MAE and 3.42% in XAUC on KuaiRec, comparing row (a) with row (h). Although different nonlinear decay strategies yield similar results in terms of XAUC, they still improve MAE. These findings indicate that the CLEM strategy improves the model’s accuracy of watch time prediction.
Figure 5. Performance comparison of different generative architectures on (a) KuaiRec and (b) CIKM16.
@@ 4.6. Performance on LTV Prediction Task (RQ5)
GR is a versatile regression framework. To explore its potential in other tasks, we conduct additional experiments on the Lifetime Value (LTV) prediction task. Following the settings  (Weng et al., 2024), we evaluate GR on Criteo-SSC [^4] https://ailab.criteo.com/criteo-sponsored-search-conversion-log-dataset/ and Kaggle [^5] https://www.kaggle.com/c/acquire-valued-shoppers-challenge, using MAE and Spearman rank correlation (Spearman’s $\rho$) as performance metrics. We consider all baseline methods compared in prior studies (Weng et al., 2024).
@@
Tab. 6 presents the results of various methods in LTV prediction. These results show that our approach significantly outperforms the previous works across ordinal and numeric performance metrics, achieving 17.66% improvement of MAE and 20.79% improvement on Spearman’s $\rho$ on Criteo-SSC, compared with the state-of-the-art method OptDist (Weng et al., 2024). Given that the compared methods are specifically optimized for the LTV task, these findings demonstrate the robustness and extensibility of GR, underscoring its potential for application to other continuous value prediction tasks within the domain of recommendation systems.
@@ 5. Conclusion
This paper proposes a novel regression paradigm Generative Regression (GR) to accurately predict watch time, which addresses two key issues associated with existing methods. First, many ordinal regression methods struggle to accurately recover durations due to discretization, with performance heavily reliant on the chosen time binning strategy. Second, while ordinal regression implicitly constrains the probability distribution along the estimation path to exhibit a decreasing trend, existing methods have not fully leveraged this property. The generative regression approach builds upon autoregressive modeling and offers a promising exploration space. We also introduce embedding mixups and curriculum learning during training to accelerate model convergence and enhance performance. Extensive offline and online experiments show that GR significantly outperforms the state-of-the-art models. Additionally, we investigate its applicability to lifetime value (LTV) prediction, where our generative regression method also surpasses the SOTA models, highlighting its potential as a novel and effective solution to general regression challenges.
Appendix A Appendices
A.1. Experimental Settings
Unless otherwise stated, we adopt the following training protocol. The hyperparameters $\lambda$ in equation Eq. (6) are deliberately set to 10. Our GR architecture comprises a two-layer Multi-Layer Perceptron (MLP) for the encoder and a single-layer Long Short-Term Memory (LSTM) network for the decoder. We employ a dropout rate of 0.3 to prevent overfitting and set the learning rate to 5e-4. As for vocabulary construction in Algorithm. 1, the decay rate $\alpha$ is set to 0.99, precision threshold $\epsilon$ is set to 1e-6, start percentile $q_{start}$ and end percentile $q_{e}nd$ are respectively set to 0.9999 and 0.90. The $\omega$ in Eq. (7) is set to 17000. In Eq. (8), the window size $n_{w}$ is set to half the length of the vocabulary. We train for 20 epochs in total with a batch size of 1024, using the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ to minimize the loss function.
Table 7. The impact of model parameters on the performance between different approaches on KuaiRec and CIKM16.
Method Parameters KuaiRec CIKM16
MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
VR 0.86M 7.634 0.504 1.039 0.641
TPM 0.86M 3.456 0.571 0.850 0.676
CREAD 0.86M 3.307 0.594 0.865 0.678
VR-large 4.34M 7.556 0.521 0.998 0.651
TPM-large 4.34M 3.432 0.577 0.843 0.682
CREAD-large 4.34M 3.293 0.599 0.857 0.680
GR (ours) 4.18M 3.196 0.614 0.815 0.691
A.2. Additional Ablation Study Results
A.2.1. Model Parameters
To ascertain that the enhancement observed in the GR model is not merely a consequence of an increase in model parameters, we conducted a comparative analysis, as presented in Tab. 7. Specifically, we standardized the parameters of the SOTA models, CREAD (Sun et al., 2024) and TPM (Lin et al., 2023), to a uniform level. The results indicate that, when evaluated with an equivalent parameter scale, the GR model consistently surpasses the previous SOTA methods across a range of metrics.
A.2.2. Hyperparameter for Vocabulary Construction
We conduct an ablation study to investigate the effects of the decay rate $\alpha$ and end percentile $q_{end}$ in Algorithm. 1 on the performance of GR. All experiments maintain a constant start percentile $q_{start}$ set at 0.9999 and the results are present in Tab. 8. Under the same end percentile $q_{end}$, models with a decay rate of 0.99 consistently outperform those with a decay rate of 0.98. Within identical settings for $\alpha$, the configurations with $q_{end}=0.9$ yield the best performance results, followed by $q_{end}=0.85$.
Figure 6. (a) Dynamic probability across different epochs for varying parameter $\omega$. The curves show the decrease in sampling rate over time. (b) The impact of $\omega$ on MAE and XAUC for the KuaiRec dataset.
A.2.3. Hyperparameter for Curriculum Learning (CL)
Here, we check the effects of the hyperparameters $\omega$ in Eq. (7) which influences the shape of descent curve to ensure a seamless transition from higher to lower values. As shown in Fig. 6(a), four distinct values of $\omega$, i.e., 4600, 8500, 13000, and 17000, are chosen to control the dynamic probability, gradually reducing it to zero at the 5th, 10th, 15th, and 20th epoch, respectively. The results are presented in Fig. 6(b). We can see that $\omega=13000$, i.e., dynamic probability approaching zero around the 15th epoch, performs better than the other settings. $\omega=4600$ (dynamic probability approaching zero around the 5th epoch) results in a premature reduction in sampling probability, leading to the worst performance. Meanwhile, the performance with $\omega=17000$, while slightly inferior to that of $\omega=13000$, still surpasses the outcomes of strategies where the probability reduces to zero too prematurely. Therefore, judiciously setting the $\omega$ parameter, and learning to predict with both ground truth and previous prediction as input can effectively resolve inconsistencies between training and inference.
Figure 7. Performance on KuaiRec dataset with different hyperparameter settings in vocabulary construction.
MAE $\downarrow$ XAUC $\uparrow$
$\alpha=0.98$ $\small q_{end}=0.95$ 3.282 0.574
$\small q_{end}=0.90$ 3.216 0.601
$\small q_{end}=0.85$ 3.227 0.589
$\small q_{end}=0.80$ 3.256 0.578
$\alpha=0.99$ $\small q_{end}=0.95$ 3.253 0.590
$\small q_{end}=0.90$ 3.196 0.614
$\small q_{end}=0.85$ 3.205 0.603
$\small q_{end}=0.80$ 3.231 0.595
Figure 8. The impact of window size $n_{w}$ on MAE and XAUC for KuaiRec dataset.
A.2.4. Hyperparameter for Embedding Mixup (EM)
Here, we check the effects of the window size $w_{n}$ in Eq. (8) which control the the fixed size of window to merge representations of tokens within it, preventing significant errors. As shown in Fig. 8, we select six percentages of the vocabulary size to determine the window size values for our ablation experiments on the Kuairec dataset. The horizontal axis represents these percentages (actual window sizes), and we also expand the window size to include weighting the entire vocabulary (denoted as ”all”). The results show that as the percentage increases, performance gradually improves, reaching its peak at 50%, and further increases in the window size lead to a slight performance decline. A smaller window size may fail to merge meaningful token representations, resulting in suboptimal performance. Conversely, a larger window size may introduce noise by incorporating less relevant tokens, which can degrade performance.

### v3

@@ 4.3. Underlying Reasons Analysis For Performance Gain (RQ2)
4.4. Vocabulary Construction Analysis (RQ3)
Figure 5. (a) Comparison of MAE on the KuaiRec dataset across videos with different watch time intervals. (b-d) The distribution comparison of predicted watch times among TPM, CREAD, and GR, compared to the Ground Truth (GT).
Table 4. Ablation study on the strategy of curriculum learning (CL) with embedding mixup (EM).
@@ 4.4. Vocabulary Construction Analysis (RQ3)
4.5. Ablation study on Curriculum Learning with Embedding Mixup (RQ4)
To systematically evaluate the proposed Curriculum Learning with Embedding Mixup (CLEM) framework, we conduct controlled ablation experiments across three dimensions:(1) component effectiveness, (2) scheduling sensitivity, and (3) nonlinear decay impact. The experimental variants a re designed as follows:
  • Component Analysis:
    • w/o CLEM: Vanilla training using direct feature projection $\bm{E}[\hat{s}_{i}^{t-1},:]\rightarrow\hat{s}_{i}^{t}$ without curriculum scheduling or mixup.
    • EM with TF: Embedding mixup with full teacher forcing (fixed sampling rate $p=1$).
    • CL w/o EM: Curriculum learning without embedding mixup regularization.
  • Decay Strategy Comparison:
    • Linear: Linear sampling rate decay $p_{t}=1-\tau t$.
    • Exponential: Exponential decay $p_{t}=e^{-\tau t}$.
  • Sampling Rate Impact:
    • Fixed-0.5: Constant sampling rate $p=0.5$.
    • Fixed-0: Pure free-running mode ($p=0$).
As shown in Tab. 4, the full CLEM framework (Row a) demonstrates significant improvements over baseline configurations. Compared to the non-curriculum variant (Row c), curriculum learning alone provides a 1.656% XAUC boost and 1.38% MAE reduction on the KuaiRec dataset. Embedding mixup contributes more substantially: disabling mixup (Row d) degrades XAUC by 4.235% and increases MAE by 4.853%, highlighting its critical regularization role. The sampling rate decay coefficients significantly impact both metrics. The proposed curriculum strategy achieves a gain of 2.65% in MAE and 3.42% in XAUC on KuaiRec, comparing row (a) with row (h). Although different nonlinear decay strategies yield similar results in terms of XAUC, they still improve MAE. These findings indicate that the CLEM strategy improves the model’s accuracy of watch time prediction.
@@ 4.6. Performance on LTV Prediction Task (RQ5)
GR is a generalized regression framework. To rigorously evaluate its cross-task generalization capability, we conduct extended experiments on the Lifetime Value (LTV) prediction task under identical experimental protocols as  (Weng et al., 2024). The evaluation employs two datasets: Criteo-SSC [^5] https://ailab.criteo.com/criteo-sponsored-search-conversion-log-dataset/ and Kaggle [^6] https://www.kaggle.com/c/acquire-valued-shoppers-challenge, with MAE and Spearman’s rank correlation (Spearman’s $\rho$) serving as performance metrics. All baseline implementations strictly adhere to the configurations documented in (Weng et al., 2024).
As shown in Tab. 5, GR achieves state-of-the-art performance with relative improvements of 17.66% in MAE and 20.79% in Spearman’s $\rho$ on Criteo-SSC over the previous best method OptDist (Weng et al., 2024). Notably, these baselines include task-specific architectures with dedicated LTV prediction modules. The consistent superiority of GR across both point estimation (MAE) and ranking correlation ($\rho$) metrics provides empirical evidence for its inherent robustness and domain-agnostic characteristics.
@@ 5. Conclusion
This paper proposes a novel regression paradigm Generative Regression (GR) to accurately predict watch time, which addresses two key issues associated with existing ordinal regression (OR) methods. First, OR struggles to accurately recover watch times due to discretization, with performance heavily reliant on the chosen time-binning strategy. Second, while OR implicitly constrains the probability distribution along the estimation path to exhibit a decreasing trend, existing methods have not fully leveraged this property. GR builds upon autoregressive modeling and offers a promising exploration space. We also introduce embedding mixups and curriculum learning during training to accelerate model convergence. Extensive online and offline experiments show that GR significantly outperforms the SOTA models. Additionally, our GR also surpasses the SOTA models in lifetime value (LTV) prediction, highlighting its potential as an effective general regression solution.